OpenStudy (anonymous):

If the Vertices Of a triangle are (sqrt7 + i sqrt3) , i(sqrt7 + i sqrt3) and sqrt7 + isqrt3 + i(sqrt7 + isqrt3) then the area of Triangle is?

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i too got it...

OpenStudy (anonymous):

thxxxx.....)))

OpenStudy (anonymous):

Two sides considered in 3D will be \[ \text{AB}=\left\{-\sqrt{3}-\sqrt{7},\sqrt{7}-\sqrt{3},0\right\}\\ \text{AC}=\left\{-\sqrt{3},\sqrt{7},0\right\}\\ AB \times AC=\{0,0,10\}\\ Area =\frac 1 2\left | AB \times AC\right|=\frac 1 2 10=5 \]

OpenStudy (anonymous):

yw